# Projected Density Matrix Sampling for Lattice Hamiltonians

### v1

@@ I Introduction
Computing the low-lying spectrum of quantum Hamiltonians is a central challenge in many-body physics and remains an active area of research across fields ranging from condensed matter to nuclear and particle physics. With the recent surge of interest in using quantum computers to solve generic quantum problems, alternative classical methods that can benchmark and cross-check quantum results are particularly valuable. While tensor networks are the main tools used today for such cross-verification [1], Monte Carlo sampling also provides a powerful framework for addressing these problems. However, its applicability is often limited by the notorious sign problem, which, in its most general form, is computationally hard [2]. The severity of the sign problem depends on several factors, including the details of the interactions, the choice of Hilbert-space basis used for Monte Carlo sampling, and the energy subspace being accessed. A variety of strategies have been explored to mitigate it in specific cases [3, 4, 5].
@@
This broad idea of extracting the low-energy spectrum using Monte Carlo techniques is well known in the lattice field theory community and is often discussed within the frame work of the generalized eigenvalue problem [6]. It is extensively employed to determine low-energy particle spectra [7, 8, 9, 10]. However, its applicability to a broad range of problems—particularly in condensed matter physics and in the Hamiltonian framework in the above form—has not yet been systematically explored. For instance, if the projection subspace coincides exactly with an energy subspace of the Hamiltonian, accurate results for the low-energy spectrum can be obtained using efficient algorithms even for arbitrarily small values of $\beta$, irrespective of whether the problem suffers from a sign problem [11]. Furthermore, by judiciously enlarging the projection subspace, the $\beta$-dependence of the extracted energies can be exponentially suppressed when accurate Monte Carlo sampling is feasible [6]. These insights, however, have not yet been effectively applied widely to physically interesting systems, like quantum spin models and qubit regularized Hamiltonians [12].
@@
Our work synthesizes several established ideas, including continuous-time path-integral formulations [13, 14], the stochastic series expansion [15, 16, 17], worm and loop algorithms [18, 19, 11], and meron-cluster techniques [20]. To illustrate the method, we apply it to two representative problems: (1) the one-dimensional Transverse Field Ising Model (tfim) in the presence of a small longitudinal magnetic field. Here, we demonstrate how our method can compute the low-lying mass spectrum of the $E_{8}$ quantum field theory that is known to emerge in this regime [21]. And, (2) the Shastry–Sutherland model (SSM) [22], a paradigmatic frustrated antiferromagnetic system with a severe sign problem but of great interest in condensed matter physics, owing to its exotic phase diagram and experimental relevance (see Ref. [23] and references therein). In this case, our goal is to investigate how well our algorithm computes the values of $e_{a}(\beta)$ on small lattices for selected projection subspaces. These thermal energies can be obtained independently using exact diagonalization for comparison. We hope that these two distinct applications of our method, along with previous developments in similar directions, help convey its broad applicability across nuclear, particle, and condensed matter physics.
@@ II Thermal Density Matrix in Continuous Time
To construct a Monte Carlo sampling procedure for the thermal density matrix, we formulate it in a continuous-time path-integral representation. [^1] While we formulate the algorithm in the path integral representation, our ideas for computing the $Z$ and $H$ matrices and using them to extract the low lying states can clearly equally well be formulated in the popular stochastic series expansion representation [15]. This is outlined in App. A. We focus on Hamiltonians that can be written in the generic form
@@ III.1 Link update
The link update modifies the space–time locations of the link operators in the configuration $\ell$]. Following the idea of the stochastic series expansion (SSE) algorithm [15], only diagonal link operators are inserted or removed during this update. A typical proposal (chosen with probability $1/2$) is either to add a new diagonal link operator or to remove an existing one. To add a link, a random spatial location (chosen uniformly among $V$ sites) and a random time slice at that location (chosen uniformly among $L_{t}$ slices) are selected, and one of the $d_{\ell}$ possible diagonal operators that can be inserted at that location is proposed. To remove a link, a random spatial location (chosen uniformly among $V$ sites) is selected, and one of the $n_{\ell}$ existing diagonal links at that location (across all time slices) is chosen uniformly and proposed for deletion. Each proposal is then accepted or rejected according to the Metropolis criterion, ensuring detailed balance.
@@ III.2 Spin update
The spin update keeps the space–time locations of the link operators $(-\varepsilon H_{\ell})$ fixed and updates the spin configuration $[s]$ in a manner consistent with the matrix elements of these operators. Several efficient methods now exist to implement changes to $[s]$ while maintaining detailed balance, including cluster updates [19], worm updates [18], and loop updates [24].
In some cases, spin configurations can be reorganized into correlated clusters. For example, in the ssm, nonzero matrix elements of the bond Hamiltonian always connect opposite spins on neighboring sites in the bra and ket states. This allows for the well known cluster representation of the spins [25]. Imposing this restriction on the link configuration in Fig. 1 generates the corresponding cluster configuration of spins, as illustrated in Fig. 2. In our work, we employ such cluster representations, together with the meron-cluster approach [20], to partially alleviate some of the sign problems associated with the ssm.
@@ IV Choice of the Projection Subspace
The choice of the projection subspace is a crucial step in the successful application of our method. This subspace determines not only the projected low-energy spectrum of $H$, but also the projected thermal energies $e_{a}(\beta)$ and the rate at which these thermal energies converge to the low-energy spectrum. We now analyze this rate of convergence quantitatively for sufficiently large values of $\beta$, where this question becomes meaningful. As we will see, the convergence is exponential in $\beta$ [6].
@@
An ideal projection subspace would be one constructed using the exact eigenstates of $H$. In that case, the second term in Eq. 21 vanishes, and the thermal energies satisfy $e_{\alpha_{i}}(\beta)={\cal E}_{\alpha_{i}}$ even for arbitrarily small $\beta$, as discussed previously. This was used recently to compute Fermi energies of free fermions using a Monte Carlo that typically suffers from a severe sign problem [11]. The next best scenario arises when perturbation theory provides a good approximation and the projection subspace of the free theory is used. In this situation, we obtain the best-case convergence with $\Delta_{a}={\cal E}_{\gamma}-{\cal E}_{\alpha_{d_{\mathrm{proj}}}}$ [6], and our method provides a non-perturbative approach that takes into account higher order corrections easily. Symmetries can also assist in constructing an effective projection subspace, since imposing symmetry constraints automatically restricts the energy eigenstates that have overlap with the chosen subspace.
@@
The associated projected low-energy level—namely, the lowest energy eigenvalue whose eigenstate has a nonzero overlap with $|\psi_{a}\rangle$—can then be determined with high accuracy by evaluating $e_{a}(\beta)$ at successively larger values of $\beta$, since the sign problem is absent. This idea is very similar to the projector Monte Carlo techniques discussed previously [26, 27].
@@ V The Transverse Field Ising Model
In the lattice theory, this prediction holds only in the scaling limit, where the lattice size $L$ is large and the magnetic perturbation $h$ is small, and the thermal perturbation is zero (tuned by the transverse field). Our goal here is to demonstrate that the pdms method can be used to confirm Eq. 24. To the best of our knowledge, previous Monte Carlo attempts to extract this ratio were not successful [28], although it has been confirmed by other numerical approaches [29, 30, 31, 32].
@@ V.1 The Hamiltonian
$H=-\sum_{i=0}^{L-1}\Big(\sigma_{i}^{z}\sigma^{z}_{i+1}-\sigma_{i}^{x}-h\,\sigma_{i}^{z}\Big),$ (25)
@@
Each $H_{\ell}$ is a nearest-neighbor bond operator that connects sites $i$ and $i+1$ such that all matrix elements of $(-\varepsilon H_{\ell})$ in the X-basis are positive. The link updates and spin updates, as discussed in Section IV are straightforward. When $h\neq 0$, the Ising symmetry is broken, but spin updates can still be constructed as explained in [33].
@@ V.4 Extraction of $E_{8}$ Mass Ratio
Our final results for the ratios $m_{2}/m_{1}$ and $m_{3}/m_{1}$ as functions of $\mu=L\,h^{8/15}$, obtained from the pdms, are shown in Fig. 5. We find that the ratios saturate to the expected $E_{8}$ value for $\mu\gtrsim 6$, within expected errors. We have verified that our results are consistent with the known universal scaling curve for these ratios as a function of $\mu$ [29].
@@ VI.1 The Hamiltonian
Using Eq. 33, we can construct the path integral and design a Monte Carlo algorithm as discussed in Sections II and III. We employ the conventional $s_{i}^{z}$ eigenbasis, labeled by $|{\uparrow}\rangle$ and $|{\downarrow}\rangle$, although other types of dimer bases could be more effective in mitigating sign problems [34]. To partially alleviate the sign problem in the $s_{i}^{z}$ basis, we work in the spin-cluster representation and include the entropy factor that arises from summing over an entire class of spin configurations during the bond updates. This helps in reducing the fluctuations while computing the $Z$ and $E$ matrices using Eqs. 17 and 18.
@@ VI.2 Projection Subspaces
for all values of $(J,J^{\prime})$. One of the central results of Ref. [22] is that, for $J^{\prime}\leq J/2$, $|\psi_{0}\rangle$ is also the ground state of $H$.
@@ VI.3 Choice of Couplings
Among the four couplings, the first case we study is $(J=0,J^{\prime}=1)$, where the ssm reduces to the 2D-unfrustrated Heisenberg quantum antiferromagnet. Here, the low-energy physics is that of a spontaneously broken continuous symmetry with a dense spectrum of low-lying excitations. Our goal in this case is to examine how well the $\mathbb{P}_{+}$ subspace reproduces the low-energy spectrum. The second coupling we consider is $(J=1,J^{\prime}=0.5)$, where the ssm lies in a gapped phase. Here, our main objective is to determine whether the energy gap can be extracted using the $\mathbb{P}_{-}$ subspace. We then investigate the couplings $(J=1,J^{\prime}=0.66)$ and $(J=0,J^{\prime}=0.75)$, which lie close to two phase transitions predicted in the ssm [23] using both the projection subspaces to learn how they perform.
@@ VI.5 Results at $L=4$
Table 6: Projected thermal energies obtained using our Monte Carlo method in the quantum Heisenberg limit for $L=8$ at various values of $\beta$. The second column gives the ground-state energy obtained using $\mathbb{P}_{(1)}$. Because this subspace is free of sign problems, the single thermal energy $e_{1}(\beta)$ can be extracted accurately even at large $\beta$. The remaining four columns give the projected thermal energies obtained from $\mathbb{P}_{+}$. Due to sign problems which sneak in from the larger projection subspace, we had to drop the projection state $|\psi_{4}\rangle$ at $\beta=5$, and used only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ at $\beta=10,15$. These results can be compared with the ground state energy of 43.103168(256)-43.103168(256) in [35], the first excited state energy of 42.90(4)-42.90(4) and the second excited state energy of 42.54(3)-42.54(3) both obtained in [36].
Figure 9: Projected thermal energies for the 2D Heisenberg square lattice, at $L=8$ as functions of inverse temperature $\beta$. The plus (purple) symbol shows the ground-state thermal energy $e_{1}(\beta)$ obtained within the single-state projection subspace $\mathbb{P}_{(1)}$, which is free of sign problems and thus yields accurate estimates even at large $\beta$. However, the results from this does not converge to the ground state (the lowest horizontal dashed line obtained in [35]). The other symbols display the projected thermal energies $e_{i}(\beta)$ ($i=1,\dots,4$) obtained from the full $\mathbb{P}_{+}$. The circle symbol gives $e_{1}(\beta)$, which is already consistent with the ground state energy at $\beta=3.0$. To avoid computational instabilities the projection state $|\psi_{4}\rangle$ was excluded from the analysis at $\beta=5$, and only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ were used at $\beta=10,15$. The higher two horizontal dashed lines are reference Monte Carlo results with their uncertainties (shaded bands) from the earlier study of Ref. [36]. These results demonstrate that (i) pdms results converge to the ground state at very low $\beta$; (ii) Comparing the blue circles with the purple plus markers, we learn that increasing the dimension of the projection subspace judiciously can considerably improve the rate of convergence.
@@
Figure 10: The energy gap between the ground state (with energy $E_{1}$) and the first excited Triplet state (with energy $E_{2}$) plotted as a function of the coupling $J^{\prime}/J$. The solid blue is obtained using perturbation theory up to $(J^{\prime}/J)^{21}$ from Ref. [37]. The colored markers without errors are ED results from a $4\times 4$ lattice. The agreement between ED and perturbative results confirms that non-perturbative effects remain small up to $J^{\prime}/J\approx 0.5$. The data from our Monte Carlo method at $J^{\prime}/J=0.5$ for $L=4$ (that also appear in Fig. 8) are shown to demonstrate the rate of convergence. We also show data from our Monte Carlo method at $J^{\prime}/J=0.5$, $L=8$ and $\beta=0.75$ and find that it agrees with the $L=4$ data within errors.

### v2

@@ I Introduction
Computing the low-lying spectrum of quantum Hamiltonians is a central challenge in many-body physics and remains an active area of research across fields ranging from condensed matter to nuclear and particle physics. With the recent surge of interest in using quantum computers to solve generic quantum problems, alternative classical methods that can benchmark and cross-check quantum results are particularly valuable. While tensor networks are the main tools used today for such cross-verification [BanulsReview23], Monte Carlo sampling also provides a powerful framework for addressing these problems. However, its applicability is often limited by the notorious sign problem, which, in its most general form, is computationally hard [TroyerWiese2005]. The severity of the sign problem depends on several factors, including the details of the interactions, the choice of Hilbert-space basis used for Monte Carlo sampling, and the energy subspace being accessed. A variety of strategies have been explored to mitigate it in specific cases [Iglovikov2015, Hangleiter2019, Pan:2022fgf].
@@
This broad idea of extracting the low-energy spectrum using Monte Carlo techniques is well known in the lattice field theory community and is often discussed within the frame work of the generalized eigenvalue problem [Blossier:2009kd]. It is extensively employed to determine low-energy particle spectra [Luscher:1990ck, DeGrand:2006zz, Esposito:2016noz, Liu:2019zoy]. Similar approaches have been proposed for constructing the reduced density matrix through a spatial bipartition for use in Monte Carlo simulations, and applied to compute entanglement entropy [PhysRevB.82.100409] and to extract entanglement level spectra in model quantum systems [PhysRevB.89.195147, mao2025sampling]. However, its applicability to a broad range of problems—particularly in condensed matter physics and in the Hamiltonian framework in the above form—has not yet been systematically explored. For instance, if the projection subspace coincides exactly with an energy subspace of the Hamiltonian, accurate results for the low-energy spectrum can be obtained using efficient algorithms even for arbitrarily small values of $\beta$, irrespective of whether the problem suffers from a sign problem [Chandrasekharan:2024iao]. Furthermore, by judiciously enlarging the projection subspace, the $\beta$-dependence of the extracted energies can be exponentially suppressed when accurate Monte Carlo sampling is feasible [Blossier:2009kd]. These insights, however, have not yet been effectively applied widely to physically interesting systems, like quantum spin models and qubit regularized Hamiltonians [Chandrasekharan:2025Cb].
@@
Our work synthesizes several established ideas, including continuous-time path-integral formulations [Beard:1996wj, RevModPhys.83.349, PhysRevB.102.094101], the stochastic series expansion [Sandvik1999_SSE, Sandvik2010_QMCReview, Sandvik2019_SSEReview], worm and loop algorithms [PhysRevLett.87.160601, Evertz:2000rk, Chandrasekharan:2024iao], and meron-cluster techniques [Chandrasekharan:1999cm]. To illustrate the method, we apply it to two representative problems: (1) the one-dimensional Transverse Field Ising Model (tfim) in the presence of a small longitudinal magnetic field. Here, we demonstrate how our method can compute the low-lying mass spectrum of the $E_{8}$ quantum field theory that is known to emerge in this regime [Zamolodchikov:1989fp]. And, (2) the Shastry–Sutherland model (SSM) [ShastrySutherland1981], a paradigmatic frustrated antiferromagnetic system with a severe sign problem but of great interest in condensed matter physics, owing to its exotic phase diagram and experimental relevance (see Ref. [Corboz2025] and references therein). In this case, our goal is to investigate how well our algorithm computes the values of $e_{a}(\beta)$ on small lattices for selected projection subspaces. These thermal energies can be obtained independently using exact diagonalization for comparison. We hope that these two distinct applications of our method, along with previous developments in similar directions, help convey its broad applicability across nuclear, particle, and condensed matter physics.
@@ II Thermal Density Matrix in Continuous Time
To construct a Monte Carlo sampling procedure for the thermal density matrix, we formulate it in a continuous-time path-integral representation. [^1] While we formulate the algorithm in the path integral representation, our ideas for computing the $Z$ and $H$ matrices and using them to extract the low lying states can clearly equally well be formulated in the popular stochastic series expansion representation [Sandvik1999_SSE]. This is outlined in App. A. We focus on Hamiltonians that can be written in the generic form
@@ III.1 Link update
The link update modifies the space–time locations of the link operators in the configuration $\ell$]. Following the idea of the stochastic series expansion (SSE) algorithm [Sandvik1999_SSE], only diagonal link operators are inserted or removed during this update. A typical proposal (chosen with probability $1/2$) is either to add a new diagonal link operator or to remove an existing one. To add a link, a random spatial location (chosen uniformly among $V$ sites) and a random time slice at that location (chosen uniformly among $L_{t}$ slices) are selected, and one of the $d_{\ell}$ possible diagonal operators that can be inserted at that location is proposed. To remove a link, a random spatial location (chosen uniformly among $V$ sites) is selected, and one of the $n_{\ell}$ existing diagonal links at that location (across all time slices) is chosen uniformly and proposed for deletion. Each proposal is then accepted or rejected according to the Metropolis criterion, ensuring detailed balance.
@@ III.2 Spin update
The spin update keeps the space–time locations of the link operators $(-\varepsilon H_{\ell})$ fixed and updates the spin configuration $[s]$ in a manner consistent with the matrix elements of these operators. Several efficient methods now exist to implement changes to $[s]$ while maintaining detailed balance, including cluster updates [Evertz:2000rk], worm updates [PhysRevLett.87.160601], and loop updates [PhysRevB.82.024407].
In some cases, spin configurations can be reorganized into correlated clusters. For example, in the ssm, nonzero matrix elements of the bond Hamiltonian always connect opposite spins on neighboring sites in the bra and ket states. This allows for the well known cluster representation of the spins [PhysRevLett.70.875]. Imposing this restriction on the link configuration in Fig. 1 generates the corresponding cluster configuration of spins, as illustrated in Fig. 2. In our work, we employ such cluster representations, together with the meron-cluster approach [Chandrasekharan:1999cm], to partially alleviate some of the sign problems associated with the ssm.
@@ IV Choice of the Projection Subspace
The choice of the projection subspace is a crucial step in the successful application of our method. This subspace determines not only the projected low-energy spectrum of $H$, but also the projected thermal energies $e_{a}(\beta)$ and the rate at which these thermal energies converge to the low-energy spectrum. We now analyze this rate of convergence quantitatively for sufficiently large values of $\beta$, where this question becomes meaningful. As we will see, the convergence is exponential in $\beta$ [Blossier:2009kd].
@@
An ideal projection subspace would be one constructed using the exact eigenstates of $H$. In that case, the second term in Eq. 21 vanishes, and the thermal energies satisfy $e_{\alpha_{i}}(\beta)={\cal E}_{\alpha_{i}}$ even for arbitrarily small $\beta$, as discussed previously. This was used recently to compute Fermi energies of free fermions using a Monte Carlo that typically suffers from a severe sign problem [Chandrasekharan:2024iao]. The next best scenario arises when perturbation theory provides a good approximation and the projection subspace of the free theory is used. In this situation, we obtain the best-case convergence with $\Delta_{a}={\cal E}_{\gamma}-{\cal E}_{\alpha_{d_{\mathrm{proj}}}}$ [Blossier:2009kd], and our method provides a non-perturbative approach that takes into account higher order corrections easily. Symmetries can also assist in constructing an effective projection subspace, since imposing symmetry constraints automatically restricts the energy eigenstates that have overlap with the chosen subspace.
@@
The associated projected low-energy level—namely, the lowest energy eigenvalue whose eigenstate has a nonzero overlap with $|\psi_{a}\rangle$—can then be determined with high accuracy by evaluating $e_{a}(\beta)$ at successively larger values of $\beta$, since the sign problem is absent. This idea is very similar to the projector Monte Carlo techniques discussed previously [PhysRevLett.95.207203, PhysRevB.82.024407].
@@ V The Transverse Field Ising Model
In the lattice theory, this prediction holds only in the scaling limit, where the lattice size $L$ is large and the magnetic perturbation $h$ is small, and the thermal perturbation is zero (tuned by the transverse field). Our goal here is to demonstrate that the pdms method can be used to confirm Eq. 24. To the best of our knowledge, previous Monte Carlo attempts to extract this ratio were not successful [Destri:1992aa], although it has been confirmed by other numerical and analytic approaches[Henkel:1989jn, Fonseca:2006au, PhysRevB.83.020407, albert2025, Jha:2024jan, litvinov2025meson]. For recent
@@ V.1 The Hamiltonian
$H=-\sum_{i=0}^{L-1}\Big(\sigma_{i}^{z}\sigma^{z}_{i+1}+\sigma_{i}^{x}+h\,\sigma_{i}^{z}\Big),$ (25)
@@
Each $H_{\ell}$ is a nearest-neighbor bond operator that connects sites $i$ and $i+1$ such that all matrix elements of $(-\varepsilon H_{\ell})$ in the X-basis are positive. The link updates and spin updates, as discussed in Section IV are straightforward. When $h\neq 0$, the Ising symmetry is broken, but spin updates can still be constructed as explained in [PhysRevE.67.046701].
@@ V.4 Extraction of $E_{8}$ Mass Ratio
Our final results for the ratios $m_{2}/m_{1}$ and $m_{3}/m_{1}$ as functions of $\mu=L\,h^{8/15}$, obtained from the pdms, are shown in Fig. 5. We find that the ratios saturate to the expected $E_{8}$ value for $\mu\gtrsim 6$, within expected errors. We have verified that our results are consistent with the known universal scaling curve for these ratios as a function of $\mu$ [Henkel:1989jn].
@@ VI.1 The Hamiltonian
Using Eq. 33, we can construct the path integral and design a Monte Carlo algorithm as discussed in Sections II and III. We employ the conventional $s_{i}^{z}$ eigenbasis, labeled by $|{\uparrow}\rangle$ and $|{\downarrow}\rangle$, although other types of dimer bases could be more effective in mitigating sign problems [PhysRevB.57.R3197]. To partially alleviate the sign problem in the $s_{i}^{z}$ basis, we work in the spin-cluster representation and include the entropy factor that arises from summing over an entire class of spin configurations during the bond updates. This helps in reducing the fluctuations while computing the $Z$ and $E$ matrices using Eqs. 17 and 18.
@@ VI.2 Projection Subspaces
for all values of $(J,J^{\prime})$. One of the central results of Ref. [ShastrySutherland1981] is that, for $J^{\prime}\leq J/2$, $|\psi_{0}\rangle$ is also the ground state of $H$.
@@ VI.3 Choice of Couplings
Among the four couplings, the first case we study is $(J=0,J^{\prime}=1)$, where the ssm reduces to the 2D-unfrustrated Heisenberg quantum antiferromagnet. Here, the low-energy physics is that of a spontaneously broken continuous symmetry with a dense spectrum of low-lying excitations. Our goal in this case is to examine how well the $\mathbb{P}_{+}$ subspace reproduces the low-energy spectrum. The second coupling we consider is $(J=1,J^{\prime}=0.5)$, where the ssm lies in a gapped phase. Here, our main objective is to determine whether the energy gap can be extracted using the $\mathbb{P}_{-}$ subspace. We then investigate the couplings $(J=1,J^{\prime}=0.66)$ and $(J=0,J^{\prime}=0.75)$, which lie close to two phase transitions predicted in the ssm [Corboz2025] using both the projection subspaces to learn how they perform.
@@ VI.5 Results at $L=4$
Table 6: Projected thermal energies obtained using our Monte Carlo method in the quantum Heisenberg limit for $L=8$ at various values of $\beta$. The second column gives the ground-state energy obtained using $\mathbb{P}_{(1)}$. Because this subspace is free of sign problems, the single thermal energy $e_{1}(\beta)$ can be extracted accurately even at large $\beta$. The remaining four columns give the projected thermal energies obtained from $\mathbb{P}_{+}$. Due to sign problems which sneak in from the larger projection subspace, we had to drop the projection state $|\psi_{4}\rangle$ at $\beta=5$, and used only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ at $\beta=10,15$. These results can be compared with the ground state energy of 43.103168(256)-43.103168(256) in [PhysRevB.56.11678], the first excited state energy of 42.90(4)-42.90(4) and the second excited state energy of 42.54(3)-42.54(3) both obtained in [PhysRevB.49.5983].
Figure 9: Projected thermal energies for the 2D Heisenberg square lattice, at $L=8$ as functions of inverse temperature $\beta$. The plus (purple) symbol shows the ground-state thermal energy $e_{1}(\beta)$ obtained within the single-state projection subspace $\mathbb{P}_{(1)}$, which is free of sign problems and thus yields accurate estimates even at large $\beta$. However, the results from this does not converge to the ground state (the lowest horizontal dashed line obtained in [PhysRevB.56.11678]). The other symbols display the projected thermal energies $e_{i}(\beta)$ ($i=1,\dots,4$) obtained from the full $\mathbb{P}_{+}$. The circle symbol gives $e_{1}(\beta)$, which is already consistent with the ground state energy at $\beta=3.0$. To avoid computational instabilities the projection state $|\psi_{4}\rangle$ was excluded from the analysis at $\beta=5$, and only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ were used at $\beta=10,15$. The higher two horizontal dashed lines are reference Monte Carlo results with their uncertainties (shaded bands) from the earlier study of Ref. [PhysRevB.49.5983]. These results demonstrate that (i) pdms results converge to the ground state at very low $\beta$; (ii) Comparing the blue circles with the purple plus markers, we learn that increasing the dimension of the projection subspace judiciously can considerably improve the rate of convergence.
@@
Figure 10: The energy gap between the ground state (with energy $E_{1}$) and the first excited Triplet state (with energy $E_{2}$) plotted as a function of the coupling $J^{\prime}/J$. The solid blue is obtained using perturbation theory up to $(J^{\prime}/J)^{21}$ from Ref. [PhysRevB.65.014408]. The colored markers without errors are ED results from a $4\times 4$ lattice. The agreement between ED and perturbative results confirms that non-perturbative effects remain small up to $J^{\prime}/J\approx 0.5$. The data from our Monte Carlo method at $J^{\prime}/J=0.5$ for $L=4$ (that also appear in Fig. 8) are shown to demonstrate the rate of convergence. We also show data from our Monte Carlo method at $J^{\prime}/J=0.5$, $L=8$ and $\beta=0.75$ and find that it agrees with the $L=4$ data within errors.